dx/dt=-4x+y+z

dy/dt=x+5y-z

dz/dt=y-3z


1
Expert's answer
2021-11-16T13:43:21-0500

he given system can be expressed as

"X'=AX" , where "X' = \begin{bmatrix}
 \frac{dx}{dt} \\[0.3em]
 \frac{dy}{dt} \\[0.3em]
 \frac{dz}{dt} \end{bmatrix},A = \begin{bmatrix}
 -1 & 4 & 2 \\[0.3em]
 4 & -1 & 2 \\[0.3em]
 0 & 0 & 6
 \end{bmatrix},X = \begin{bmatrix}
 x \\[0.3em]
 y \\[0.3em]
 z \end{bmatrix}"

Find the eigenvalues and eigenvectors of the coefficient matrix "A".

Characteristic equation of the matrix "A" is given by

"det(A-\lambda I)=0"

"\Rightarrow \left | 
 \begin{matrix}
 \lambda +1 & 4 & 2 \\
 4 & \lambda +1 & 2 \\
 0 & 0 & 6-\lambda 
 \end{matrix}
\ \right |=0"

"\Rightarrow -\lambda ^3+4\lambda ^2+27\lambda -90=0"

"\Rightarrow (\lambda +5)(\lambda -3)(\lambda -6)=0"

"\Rightarrow \lambda =-5,3,6"

So, the eigenvalues of the matrix "A" are: "\Rightarrow \lambda =-5,3,6"

Now let us find the eigenvectors corresponding to the eigenvalues.

Eigenvector corresponding to the eigenvalue "\lambda =-5" :

Let "v_{1}" be the eigenvector corresponding to the eigenvalue "\lambda=-5"

Then, "(A+5*I)v=0"

"\Rightarrow \begin{bmatrix}
 4 & 4 & 2 \\[0.3em]
 4 & 4 & 2 \\[0.3em]
 0 & 0 & 11
 \end{bmatrix}v=0"

Solve the homogeneous system by Gaussian elimination.

"\Rightarrow \begin{bmatrix}
 2 & 2 & 1 \\[0.3em]
 0 & 0 & 0 \\[0.3em]
 0 & 0 & 1
 \end{bmatrix}v=0"

"\Rightarrow x_{1}+x_{2}=0,x_{3}=0"

"\Rightarrow v_{1}=\begin{bmatrix}
 -1 \\[0.3em]
 1 \\[0.3em]
 0 
 \end{bmatrix}"

Similarly, the eigenvectors corresponding to the eigenvalues "\lambda=3,6" are

"v_{2}=\begin{bmatrix}
 1 \\[0.3em]
 1 \\[0.3em]
 0 
 \end{bmatrix},v_{3}=\begin{bmatrix}
 2 \\[0.3em]
 2 \\[0.3em]
 3 
 \end{bmatrix}"

Therefore, the general solution of the given system is

"X(t)= c_{1}e^{-5t}\begin{bmatrix}
 -1 \\[0.3em]
 1 \\[0.3em]
 0 
 \end{bmatrix}+c_{2}e^{3t}\begin{bmatrix}
 1 \\[0.3em]
 1 \\[0.3em]
 0 
 \end{bmatrix}+c_{3}e^{6t}\begin{bmatrix}
 2 \\[0.3em]
 2 \\[0.3em]
 3 
 \end{bmatrix}"

That is, "\begin{bmatrix}
 x(t) \\[0.3em]
 y(t) \\[0.3em]
 z(t) 
 \end{bmatrix}= \begin{bmatrix}
 - c_{1}e^{-5t} +c_{2} e^{3t}+2c_{3}e^{6t} \\[0.3em]
 c_{1}e^{-5t} +c_{2} e^{3t}+2c_{3}e^{6t} \\[0.3em]
 3c_{3}e^{6t} 
 \end{bmatrix}"